OpenStudy (anonymous):

Janine made a cylindrical vase in which the sum of the lateral area and area of one base was about 3000 square centimeters. The vase had a height of 50 centimeters. Find the radius of the vase. Explain the method you would use to find the radius.

OpenStudy (anonymous):

?

OpenStudy (kropot72):

lateral area of cylinder = circumference * height\[A _{l}=2\times \pi \times r \times h\ ...........(1)\] area of base is given by\[A _{b}=\pi \times r ^{2}\ ...........(2)\] \[A _{l}+A _{b}=3000=(2\times \pi \times r \times 50)+(\pi \times r ^{2})\ ........(3)\] You can form a quadratic equation from (3) and one of the solutions is the radius.

OpenStudy (anonymous):

is this the answer because i don't understand at all..lol

OpenStudy (kropot72):

The quadratic equation is as follows: \[\pi r ^{2}+100\pi r-3000=0\] You need to solve this to find r the radius of the vase.

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

8.78 cm

OpenStudy (anonymous):

@kropot72

OpenStudy (anonymous):

is my answer right 8.78 cm

OpenStudy (kropot72):

How did you get that answer?